# Mieko Ochi

### v2

@@ Introduction
Hematoxylin and eosin–stained mitotic figure (MF) counts are essential for tumor evaluation, serving both as standalone and component grades in malignancy assessment (1). Mitotic figures are broadly classified into typical and atypical variants, with atypical forms—characterized by dysregulated chromatin aggregation and reflecting genomic instabilities such as chromosomal instability and aneuploidy—demonstrating independent prognostic value in cancers like breast carcinoma (2, 3). However, manual enumeration and discrimination of MF variants are time-consuming and subject to substantial inter-observer variability.
@@ 0.2 LoRA for the Foundation Models and Ensemble Learning
We fine-tuned ImageNet-pretrained ConvNeXt V2 (base, fcmae_ft_in22k_in1k) with a binary classification head. For PFMs, based on the results reported by Banerjee et al. from LoRA fine-tuning of foundation models on the AMi-Br dataset (11), we selected three PFMs — UNI (15), Virchow (16), and Virchow2 (17) — for our own LoRA fine-tuning experiments. Furthermore, following Haotian et al.’s method (18) of applying LoRA to the query (Q) and value (V) projection matrices in the multi-head self-attention (MHSA) module, we introduced two low-rank matrices $A_{Q}\in\mathbb{R}^{d\times r}$ and $B_{Q}\in\mathbb{R}^{r\times k}$, such that the update to the frozen query weight $W_{Q}\in\mathbb{R}^{d\times k}$ can be factorized as
@@
Only $A_{Q}$ and $B_{Q}$ are learned during fine-tuning, while the original weight matrix $W_{0}$ remains fixed.
@@
Here, $C$ denotes the number of classes and $D_{\mathrm{val}}^{c}$ denotes the set of validation samples with label $c$. By equally weighting the accuracy of each class, we mitigate performance degradation on minority classes.
@@ Discussion
We based our approach on the first-place solution by Haotian et al (18) in the Pap Smear Cell Classification Challenge, combining parameter–efficient LoRA fine-tuning of individual PFMs with a subsequent ensemble learning stage. Banerjee et al. (19) demonstrated that, among the diverse PFMs evaluated on the AMi-Br dataset (one of the challenge’s benchmarks) (11), Virchow2, UNI, and Virchow achieved the high balanced accuracies. Accordingly, we selected these three models for per-model adaptation. Each PFM has been pretrained on large collections of human histopathology images acquired under different conditions (institutions, countries, staining protocols, scanners, etc.), resulting in complementary attention patterns over cellular structures. By ensembling their predictions, we exploit these diverse inductive biases to arrive at more accurate final diagnoses. In addition, we incorporated ConvNeXt V2, a convolutional neural network (CNN) that achieved state-of-the-art performance across multiple vision benchmarks, including ImageNet recognition, in 2023, into the ensemble. Owing to its architectural design, CNNs extract features within local receptive fields of convolutional kernels, thereby capturing local patterns. In contrast, Vision Transformers extract features in parallel across the entire image, resulting in global receptive fields and holistic representations (20, 21). Given that the target regions of interest in our task—mitotic figures—are confined to small local areas, we hypothesized that CNNs would provide superior accuracy.

### v3

@@ Introduction
Hematoxylin and eosin–stained mitotic figure (MF) counts are essential for tumor evaluation, serving both as standalone and component grades in malignancy assessment (1). Mitotic figures are broadly classified into typical and atypical variants, with atypical forms—characterized by dysregulated chromatin aggregation and reflecting genomic instabilities such as chromosomal instability and aneuploidy—demonstrating independent prognostic value in cancers like breast carcinoma (2, 3). However, manual enumeration and discrimination of MF variants are time-consuming and subject to substantial inter‐observer variability.
@@ 0.2 LoRA for the Foundation Models and Ensemble Learning
We fine-tuned ImageNet-pretrained ConvNeXt V2 (base, fcmae_ft_in22k_in1k) with a binary classification head. For PFMs, based on the results reported by Banerjee et al. from LoRA fine-tuning of foundation models on the AMi-Br dataset (11), we selected three PFMs — UNI (15), Virchow (16), and Virchow2 (17) — for our own LoRA fine-tuning experiments. Furthermore, following Haotian et al.’s method (18) of applying LoRA to the query (Q) and value (V) projection matrices in the multi-head self-attention (MHSA) module, we introduced two low‐rank matrices $A_{Q}\in\mathbb{R}^{d\times r}$ and $B_{Q}\in\mathbb{R}^{r\times k}$, such that the update to the frozen query weight $W_{Q}\in\mathbb{R}^{d\times k}$ can be factorized as
@@
Only $A_{Q}$ and $B_{Q}$ are learned during fine‐tuning, while the original weight matrix $W_{0}$ remains fixed.
@@
Here, $C$ denotes the number of classes and $D_{\mathrm{val}}^{c}$ denotes the set of validation samples with label $c$. By equally weighting the accuracy of each class, we mitigate performance degradation on minority classes. Equation 4 defines our proposed objective for the ensemble stage (i.e., maximizing balanced accuracy) under simplex constraints on the weights. The ensemble weights were not updated jointly with base models. Instead, all individual models were fully trained first, and the weights were then optimized post hoc on the validation set using AutoGluon’s weighted ensemble (Caruana-style Ensemble Selection) (19). Concretely, we employ a greedy forward selection with replacement over the candidate model library, iteratively adding the model that yields the largest increase in balanced accuracy on out-of-fold/validation predictions, and normalizing the nonnegative weights to sum to 1. This procedure is implemented via AutoGluon Tabular (v1.2) and corresponds to its default weighted-ensemble meta-learner.
@@ Discussion
We based our approach on the first‐place solution by Haotian et al. (18) in the Pap Smear Cell Classification Challenge, combining parameter–efficient LoRA fine-tuning of individual PFMs with a subsequent ensemble learning stage. Banerjee et al. (20) demonstrated that, among the diverse PFMs evaluated on the AMi-Br dataset (one of the challenge’s benchmarks) (11), Virchow2, UNI, and Virchow achieved the high balanced accuracies. Accordingly, we selected these three models for per-model adaptation. Each PFM has been pretrained on large collections of human histopathology images acquired under different conditions (institutions, countries, staining protocols, scanners, etc.), resulting in complementary attention patterns over cellular structures. By ensembling their predictions, we exploit these diverse inductive biases to arrive at more accurate final diagnoses. In addition, we incorporated ConvNeXt V2, a convolutional neural network (CNN) that achieved state-of-the-art performance across multiple vision benchmarks, including ImageNet recognition, in 2023, into the ensemble. Owing to its architectural design, CNNs extract features within local receptive fields of convolutional kernels, thereby capturing local patterns. In contrast, Vision Transformers extract features in parallel across the entire image, resulting in global receptive fields and holistic representations (21, 22). Given that the target regions of interest in our task—mitotic figures—are confined to small local areas, we hypothesized that CNNs would provide superior accuracy.